# Quantum entanglement in phase space

### v1

@@ Abstract
Most criteria designed to detect entanglement in continuous variable (CV) systems rely on quadrature measurements. A primary example is the Simon [1] and Duan et al. [2] criteria, which state that for all separable states it holds
@@
where $\hat{x}_{k}$ and $\hat{p}_{k}$ are respectively the position and momentum quadratures of system $k\in\{A,B\}$ satisfying $[\hat{x}_{k},\hat{p}_{l}]=2i\delta_{kl}$. Observing a violation of Eq. (1) thus implies entanglement between $A$ and $B$. Moreover, a simple generalization of this condition has been shown to be necessary and sufficient for detecting entanglement in Gaussian states [2].
@@
Known examples in this direction are the Bell inequality by Banaszek and Wódkiewicz [8], which allows for the detection of nonlocality from parity measurements at only four points in phase-space, and the Einstein-Podolsky-Rosen (EPR) criterion by Walschaers and Treps [9], which allows for the detection of steering when a state with negative Wigner function is remotely prepared. These criteria, however, are intrinsically much more demanding, since revealing nonlocality and steering requires much lower levels of noise compared to entanglement. Recently, some works pointed out that performing a balanced beam-splitter transformation on $\rho_{AB}$ allows for revealing entanglement properties from the reduced Wigner function of one of the transformed modes [10, 11]. These results suggest the possibility of deriving entanglement from a few measurements in phase-space, without the need of full tomography.
@@
where $[x^{\prime}=x^{\prime}(x,p),p^{\prime}=p^{\prime}(x,p)]$ is an arbitrary linear transformation of the coordinate system for $B$, see Eq. (7), and $R$ is an arbitrary region in phase space. We provide a method to derive the optimal transformation and show that known criteria [1, 2, 10, 11] can be understood as specific choices. In particular, we prove that for Gaussian states criterion (I) is necessary and sufficient, as it coincides with Duan-Simon criteria, and that criterion (III) can be expressed as the measurement of the reduced Wigner function at a single point in the phase space of a transformed basis. We then show examples demonstrating how our criteria successfully detect entanglement in a variety of routinely prepared Gaussian and non-Gaussian CV states, such as two-mode squeezed states, Werner states and entangled cat states.
@@
Wigner function formalism.– Consider a system of $N$ bosonic modes, each described by creation and annihilation operators $\hat{a}_{k}^{\dagger}$ and $\hat{a}_{k}$, which satisfy the commutation relation $\left[\hat{a}_{k},\hat{a}_{k^{\prime}}^{\dagger}\right]=\delta_{kk^{\prime}}$ for $k,k^{\prime}\in\{1,\dots,N\}$. For each mode, it is natural to introduce position and momentum operators $\hat{x}_{k}=\hat{a}_{k}^{\dagger}+\hat{a}_{k}$ and $\hat{p}_{k}=i\left(\hat{a}_{k}^{\dagger}-\hat{a}_{k}\right)$, satisfying $[\hat{x}_{k},\hat{p}_{k^{\prime}}]=2i\delta_{kk^{\prime}}$. We arrange these in the vector $\hat{\bm{\xi}}=\left(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N}\right)^{\top}$, which is associated to the vector of phase space coordinates $\bm{\xi}\equiv\left(x_{1},p_{1},\ldots,x_{N},p_{N}\right)^{\top}$. A convenient representation of the state $\rho$ of a CV system is given by the Wigner function
@@
The rest of this work will deal with a bipartite, i.e. $N=2$, system described by the density matrix $\rho_{AB}$ or, equivalently, by the joint Wigner function $W_{AB}(x_{A},p_{A},x_{B},p_{B})$. Our goal is to detect entanglement between the $A$ and $B$ subsystems based on this function. Note that such a result is also useful towards revealing multipartite entanglement, as one can test inseparability across all possible bipartitions of the $N>2$ subsystems.
@@
A determinant $\Delta=+1$ indicates a canonical transformation preserving the commutator $[x^{\prime},p^{\prime}]=[x,p]=i\hbar$. Such a transformation of the quadratures $(x,p)$ is often referred to as a symplectic transformation, where $(x,p)$ and $(x^{\prime},p^{\prime})$ are related through a unitary transformation, $x^{\prime}=U^{\dagger}xU,p^{\prime}=U^{\dagger}pU$, as their spectra are preserved up to a constant displacement $(x_{0},p_{0})$ [18]. To transition from the Heisenberg picture to the Schrödinger picture, we move from a unitary transformation of the operator to a unitary transformation of the quantum state. This unitary transformation affects the Wigner function as $W_{U\rho U^{\dagger}}(x^{\prime},p^{\prime})=W_{\rho}\left(x,p\right)$. Typical unitary transformations are displacement, rotation and single mode squeezing.
@@
which is still a valid Wigner function.
@@
For example, taking the axis of the mirror reflection as the $x$-axis, such a transformation has already been used to generalize the well-known Peres-Horodecki separability criterion [20, 21] to continuous variable systems [1], since the transpose operation of the density matrix $\hat{\rho}\rightarrow\hat{\rho}^{T}$ is equivalent to the mirror reflection $W(x,p)\rightarrow W(x,-p)$ in phase space. The core concept of the Simon criteria is to have the partially transposed state violate the uncertainty relation. As further explained in the SM Sec.I [22], we demonstrate that for a single-mode Gaussian state, such a violation leads to a state purity exceeding unity. Additionally, we provide an intuitive phase-space analogous to the Simon criterion in the SM Sec.I [22].
@@
Apart from mirror reflection along the $x$-axis, the criteria (I)-(III) we present are more general, since they consider any linear transformation $(x,p)\rightarrow(x^{\prime},p^{\prime})$ of the coordinate system that follows Eq. (7). In addition, our criteria do not require full information of the state. We now give a proof of these criteria.
@@
Proof of criterion (I) – Consider an arbitrary transformation $(x^{\prime},p^{\prime})$ which can be determined by the real parameters $a,b,c,d,x_{0},p_{0}$. Based on these parameters, we then construct a new single-mode transformation
$h:\left(\begin{array}[]{c}x\\ p\end{array}\right)\rightarrow\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\left(\begin{array}[]{c}x+\frac{X}{\cos\theta}\\ p+\frac{P}{\cos\theta}\end{array}\right)+\left(\begin{array}[]{c}\sin\theta x_{0}\\ \sin\theta p_{0}\end{array}\right)$ (10)
@@
where $X,P$ are arbitrary constant coordinates. Assume $\rho_{AB}$ is a separable state with a Wigner function given by Eq. (2). The left-hand side of criterion (I) can be rewritten as
$\iint_{-\infty}^{\infty}\sum_{\lambda}p(\lambda)W_{A}^{\lambda}(\cos\theta x,\cos\theta p)$ (11)
×$\times W_{h\left(B\right)}^{\lambda}\left(\sin\theta x-\frac{X}{\cos\theta},\sin\theta p-\frac{P}{\cos\theta}\right)\text{d}x\text{d}p$
$=$ $\iint_{-\infty}^{\infty}\sum_{\lambda}p(\lambda)W_{A}^{\lambda}\left(\cos\theta x+\sin\theta X,\cos\theta p+\sin\theta P\right)$
×$\times W_{h\left(B\right)}^{\lambda}\left(\sin\theta x-\cos\theta X,\sin\theta p-\cos\theta P\right)\text{d}x\text{d}p\;.$
@@
This can be interpreted as follows: Perform a beam-splitter transformation $\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{array}\right)$ to the two modes $A$ and $h(B)$, and denote the resulting modes as $A^{\prime}:(x,p)$ and $B^{\prime}:(X,P)$. Notice that Eq. (11) is exactly the Wigner function of mode $B^{\prime}$, with mode $A^{\prime}$ integrated out. Therefore Eq. (11) should be upper-bounded by $1/(2\pi)$. Detailed proof can be found in Sec.II of SM [22]. We further show that criterion (I) is necessary and sufficient for Gaussian states in the Sec.III of SM [22]. $\Box$
@@
×$\quad\times\sqrt{\left(\iint_{R}\left[W_{f\left(\rho_{B}\right)}^{\lambda}(x\sin\theta,p\sin\theta)\right]^{2}\text{d}x\text{d}p\right)}$

### v2

@@ Abstract
Introduction
Most criteria designed to detect entanglement in experiments with continuous variable (CV) systems rely on quadrature measurements. A primary example is the Simon [1] and Duan et al. [2] criteria, which state that for all separable states
@@ Introduction
where $\hat{x}_{k}$ and $\hat{p}_{k}$ are respectively the position and momentum quadratures of system $k\in\{A,B\}$ satisfying $\hat{x}_{k},\hat{p}_{l}]=2i\delta_{kl}$. Observing a violation of Eq. (1) thus implies entanglement between $A$ and $B$. Moreover, a simple generalization of this condition has been shown to be necessary and sufficient for detecting entanglement in Gaussian states [2].
@@
Known examples in this direction are the Bell inequality by Banaszek and Wódkiewicz [8, 9], which allows for the detection of nonlocality from parity measurements at only four points in phase-space, and the Einstein-Podolsky-Rosen (EPR) criterion by Walschaers and Treps [10], which allows for the detection of steering when a state with negative Wigner function is remotely prepared. These criteria, however, are intrinsically much more demanding, since revealing nonlocality and steering requires much lower levels of noise compared to entanglement. Recently, some works pointed out that performing a balanced beam-splitter transformation on $\rho_{AB}$ allows for revealing entanglement properties from the reduced Wigner function of one of the transformed modes [11, 12]. These results suggest the possibility of deriving entanglement from a few measurements in phase-space, without the need of full tomography.
@@
where $x^{\prime}=x^{\prime}(x,p),p^{\prime}=p^{\prime}(x,p)]$ is an arbitrary linear transformation of the coordinate system for $B$, see Eq. (8) and Fig. 1, and $R$ is an arbitrary region in phase space. We provide a method to derive the optimal transformation and show that known criteria [1, 2, 11, 12, 13] can be understood as specific choices. In particular, we prove that for Gaussian states criterion (I) is necessary and sufficient, as it coincides with Duan-Simon criteria, and that criterion (III) can be expressed as the measurement of the reduced Wigner function at a single point in the phase space of a transformed basis. We then show examples demonstrating how our criteria successfully detect entanglement in a variety of routinely prepared Gaussian and non-Gaussian CV states, such as two-mode squeezed states, Werner states and entangled cat states.
Interestingly, we prove that Criterion (I) has a close connection with the entanglement negativity and can thus be used as a bound on entanglement monotones. This originates from a connection between Criterion (I) and the Peres-Horodecki separability criterion [14, 15], which states that all separable states must have positive partial transpose (PPT). This follows from the fact that $\rho_{AB}^{T_{B}}=\sum_{\lambda}p(\lambda)\rho_{A}^{\lambda}(\rho_{B}^{\lambda})^{T}$ is a valid state with non-negative eigenvalues. A violation of this condition indicates entanglement. Such criterion is particularly powerful in CV systems, where bound entanglement is a rare phenomenon [16]. For entangled states that are not PPT, the entanglement negativity is a useful entanglement monotone defined via the negative eigenvalues of $\rho_{AB}^{T_{B}}$ [17, 18, 19, 20], specifically, $\mathcal{N}(\rho)=(\|\rho^{T_{B}}\|_{\text{tr}}-1)/2$. In the SM Sec. I [21], we show that the violation $v$ of criterion (I), quantified as the difference between both sides, provides a lower bound for entanglement negativity $\mathcal{N}(\rho)\geq\pi v$. For Schmidt-correlated states, both the PPT criterion [22] and criterion (I) are sufficient and necessary conditions. Moreover, the optimal $\pi v$ can even accurately provide the exact value of the entanglement negativity.
Similarly, criterion (II) exhibits a fundamental connection with the computable cross norm or realignment (CCNR) criterion [23, 24, 25, 26], which complements the PPT criterion. The left-hand side of criterion (II) can be expressed as
$\text{l.h.s}=\frac{1}{2\pi|\sin 2\theta|}\iint_{R}\left|\left\langle\frac{|\cos\theta|}{\sqrt{\pi}}D(x\cos\theta,p\cos\theta)\hat{\Pi}D^{\dagger}(x\cos\theta,p\cos\theta)\otimes\frac{|\sin\theta|}{\sqrt{\pi}}D(x^{\prime}\sin\theta,p^{\prime}\sin\theta)\hat{\Pi}D^{\dagger}(x^{\prime}\sin\theta,p^{\prime}\sin\theta)\right\rangle\right|\text{d}x\text{d}p,$ (3)
where $D\hat{\Pi}D^{\dagger}$ represents the displaced parity measurement detailed in the following section. The operators $P_{A}(x,p)=\frac{|\cos\theta|}{\sqrt{\pi}}D(x\cos\theta,p\cos\theta)\hat{\Pi}D^{\dagger}(x\cos\theta,p\cos\theta)$ and $P_{B}(x^{\prime},p^{\prime})=\frac{|\sin\theta|}{\sqrt{\pi}}D(x^{\prime}\sin\theta,p^{\prime}\sin\theta)\hat{\Pi}D^{\dagger}(x^{\prime}\sin\theta,p^{\prime}\sin\theta)$ are specifically chosen to be normalized such that $\operatorname{tr}(P^{\dagger}(x_{1},p_{1})P(x_{2},p_{2}))=\delta(x_{1}-x_{2})\delta(p_{1}-p_{2})$. Consequently, the integrand represents a Bloch decomposition of the quantum state [27]. Following the analysis presented in Ref. [28] and using properties of the trace norm, we obtain $\text{l.h.s.}\leq\frac{\|\mathcal{R}(\rho)\|_{\operatorname{tr}}}{2\pi|\sin 2\theta|}$, where $\mathcal{R}(\rho)$ denotes the realigned density matrix in the CCNR criterion.
Results
@@
Wigner function formalism.– Consider a system of $N$ bosonic modes, each described by creation and annihilation operators $\hat{a}_{k}^{\dagger}$ and $\hat{a}_{k}$, which satisfy the commutation relation $\left[\hat{a}_{k},\hat{a}_{k^{\prime}}^{\dagger}\right]=\delta_{kk^{\prime}}$ for $k,k^{\prime}\in\{1,\dots,N\}$. For each mode, it is natural to introduce position and momentum operators $\hat{x}_{k}=\hat{a}_{k}^{\dagger}+\hat{a}_{k}$ and $\hat{p}_{k}=i\left(\hat{a}_{k}^{\dagger}-\hat{a}_{k}\right)$, satisfying $\hat{x}_{k},\hat{p}_{k^{\prime}}]=2i\delta_{kk^{\prime}}$. We arrange these in the vector $\hat{\bm{\xi}}=\left(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N}\right)^{\top}$, which is associated to the vector of phase space coordinates $\bm{\xi}\equiv\left(x_{1},p_{1},\ldots,x_{N},p_{N}\right)^{\top}$. A convenient representation of the state $\rho$ of a CV system is given by the Wigner function
@@
The rest of this work will deal with a bipartite, i.e. $N=2$, system described by the density matrix $\rho_{AB}$ or, equivalently, by the joint Wigner function $W_{AB}(x_{A},p_{A},x_{B},p_{B})$. Our goal is to detect entanglement between the $A$ and $B$ subsystems based on this function. Crucially, as illustrated in Fig. 1, our criteria do not require characterizing the entire four-dimensional Wigner function. Instead, they rely on specific two-dimensional slices where the variables $(x_{B},p_{B})$ are constrained to be functions of $(x_{A},p_{A})$, reducing the measurement complexity from four independent variables to just two. (For notational simplicity, we then drop the subscript $A$.) Note that such a result is also useful towards revealing multipartite entanglement, as one can test inseparability across all possible bipartitions of the $N>2$ subsystems.
@@
A determinant $\Delta=+1$ indicates a canonical transformation preserving the commutator $x^{\prime},p^{\prime}]=[x,p]=i\hbar$. Such a transformation of the quadratures $(x,p)$ is often referred to as a symplectic transformation, where $(x,p)$ and $(x^{\prime},p^{\prime})$ are related through a unitary transformation, $x^{\prime}=U^{\dagger}xU,p^{\prime}=U^{\dagger}pU$, as their spectra are preserved up to a constant displacement $(x_{0},p_{0})$ [35]. To transition from the Heisenberg picture to the Schrödinger picture, we move from a unitary transformation of the operator to a unitary transformation of the quantum state. This unitary transformation affects the Wigner function as $W_{U\rho U^{\dagger}}(x^{\prime},p^{\prime})=W_{\rho}\left(x,p\right)$. Typical unitary transformations are displacement, rotation and single mode squeezing.
@@
which is still a valid Wigner function. In the SM Sec. II [21], we discuss the choices of the determinant $\Delta=\pm 1$ for different criteria. Criterion (I) is nontrivial when $\Delta=-1$ and Criterion (III) is nontrivial when $\Delta=1$.
@@
For example, taking the axis of the mirror reflection as the $x$-axis, such a transformation has already been used to generalize the Peres-Horodecki separability criterion [14, 15] to continuous variable systems [1], since the transpose operation of the density matrix $\hat{\rho}\rightarrow\hat{\rho}^{T}$ is equivalent to the mirror reflection $W(x,p)\rightarrow W(x,-p)$ in phase space. The Simon criterion detects entanglement by showing that the partially transposed state violates the uncertainty relation that separable states must satisfy [1]. For two-mode Gaussian states, when the Simon criterion is violated and the partially transposed state is sent through a beam splitter, one output mode exhibits an extremely localized Wigner function that would correspond to an unphysical purity exceeding unity, as detailed in SM Sec. III [21]. In contrast, our criterion (I) employs a fundamentally different approach: rather than testing uncertainty relation violations, it detects when the partially transposed state violates the upper bound of the Wigner function that all physical states must satisfy, in closer analogy to the original PPT criterion. This distinction, which we illustrate through a phase-space interpretation in SM Sec. III [21], explains why criterion (I) can outperform the Simon criterion for certain non-Gaussian states, as demonstrated in Fig. 8. Furthermore, for Gaussian states, the Simon criterion essentially utilizes the complete covariance matrix information, which fully determines the shape of the joint Wigner function (the first-order moments only affect the position of the Gaussian distribution). In contrast, criterion (I) relies solely on partial information from a specific phase-space slice, requiring less data than that needed for full reconstruction of the joint Wigner function, while still enabling reliable entanglement detection.
@@
Apart from mirror reflection along the $x$-axis, the criteria (I)-(III) we present are more general, since they consider any linear transformation $(x,p)\rightarrow(x^{\prime},p^{\prime})$ of the coordinate system that follows Eq. (8). In addition, our criteria do not require full information of the state. We give a proof of these criteria in the Methods section. We further show that criterion (I) is necessary and sufficient for Gaussian states in the Sec. V of SM [21].
Testing criterion (III) can be performed in two ways. The first is to measure it in its original form, which is equivalent to performing single-mode tomography to obtain a full slice of the two-mode Wigner function. Alternatively, if the transformation $(x^{\prime},p^{\prime})$ corresponds to a unitary transformations acting on $\rho_{B}$, a second, potentially weaker method can be applied, requiring to measure only one point of the reduced Wigner function. Specifically, let us consider a beam-splitter transformation of the modes, $\hat{a}_{\pm}=(\hat{a}_{A}\pm\hat{a}_{B})/\sqrt{2}$. Using Eq. (15) with $\theta=\pi/4$ and $\left(x^{\prime}=\sqrt{2}x_{+}-x,p^{\prime}=\sqrt{2}p_{+}-p\right)$, we arrive at the special case $\iint_{-\infty}^{\infty}W_{AB}\left(x,p,x^{\prime},p^{\prime}\right)\text{d}x\text{d}p=\frac{1}{2}W_{+}\left(x_{+},p_{+}\right)$, which is half the Wigner function of one of the two modes resulting from the beam-splitter transformation. This means that, to violate criterion (III) it is enough to find a single point in the $(x_{+},p_{+})$ space where $W_{+}$ is negative [11, 12], see SM Sec. VI [21].
Applications and comparison to known criteria.– In the following, we illustrate how criteria (I,II,III) are applied to several experimentally relevant scenarios. For a comparison between our entanglement criteria and nonlocality/steering criteria based on the Wigner function see SM Sec. VII [21].
In the simplest case of TMSV states, i.e. $\eta=1,r=0$, we find that the linear transformation $(x^{\prime}=x,p^{\prime}=-p)$ and $\theta=\pi/4$ are optimal and allow us to detect entanglement with both criteria (I,II) as long as $s>0$ (See SM Sec. VIII [21]). This result coincides with the conclusion obtained from Duan-Simon and PPT criteria, which are known to be necessary and sufficient [1, 2, 14, 39].
In the most general case of a TMST state, i.e. $1>\eta>0,r>0$, Duan-Simon and PPT criteria are still necessary and sufficient conditions for entanglement [1, 2, 14, 39]. Remarkably, we can prove that this is also true for our criterion (I), see SM Sec. V [21]. On the other hand, while our criterion (II) is still useful, the bound it provides is not tight. In fact, for Gaussian states criterion (II) is a specific case of criterion (I) with $\theta=\pi/4$ and $R=\mathbb{R}^{2}$, and therefore (I) is strictly stronger than (II). In Fig. 2, we illustrate this with a concrete example. The shaded region indicates TMST states that are entangled. These states are all detected by Simon’s criterion, which is violated above the black solid line, since it is necessary and sufficient for Gaussian states [1]. The same conclusion can be obtained from our criterion (I). In addition, our criterion (II) detects only a fraction of all TMST entangled states, as indicated by the regions above the dashed coloured lines. Note that these depend on the squeezing $s$, while the bound obtained from criterion (I) is independent of $s$. Contrary to (I,II), entanglement criteria in Refs. [11, 12] fail to detect entanglement for any TMST states, more precisely, any Gaussian state, since they only work for states with Wigner negativity.
[…]
Let us start with Werner state $\rho(\Phi^{+})$, whose Wigner distribution $W_{AB}$ under the linear transformation $x^{\prime}=x,p^{\prime}=-p$ (or equivalently, $a=1,b=c=0,d=-1,x_{0}=p_{0}=0$ in Eq. 8) is non-negative for any $\epsilon$ (See SM Sec. VIII [21]). Considering the integration to be over the entire space $\mathbb{R}^{2}$ and $\theta=\pi/4$, we obtain for criterion (II)
which is violated for $\epsilon>1/3$. Since $\theta=\pi/4$ represents a balanced setting, we have moved the constant it introduces to the other side of the inequality. The same conclusion is obtained from criterion (I). This result coincides with the known tight bound for entanglement of $\rho(\Phi^{+})$ [39].
For Werner state $\rho(\Psi^{+})$, we find that the Wigner function $W_{AB}(x,p,-x,-p)$ is negative over the entire phase space at $\epsilon=1$ (See SM Sec. VIII [21]). We thus calculate criterion (III) and obtain
Figure 3: Entanglement detection for dephased cat states Eq. (13). (a): Slice of the joint Wigner function defined by $W_{AB}(x,p,x,-p)$, for $\gamma=2$ and $\epsilon=1$. Yellow dashed circles indicate the smallest region $R$ needed to violate criterion (II). (b): Amount of violation for criterion (II) with $\theta=\pi/4$, i.e. $\iint_{-\infty}^{\infty}|W_{AB}(x,p,x,-p)|\text{d}x\text{d}p-1/(4\pi)$, and for PPT criterion, i.e. $|\lambda_{\min}(\rho^{T_{B}})|$, as a function of $\epsilon$ and for three values of $\gamma$. Entanglement can be revealed by both criteria whenever $\gamma,\epsilon>0$.
Dephased two-mode cat states.– To illustrate the performance of our criteria on non-Gaussian states with many excitations, we consider states of the form
As illustrated in Fig. 3(a), the joint Wigner distribution under the transformation $(x^{\prime}=x,p^{\prime}=-p)$ and $\theta=\pi/4$, which we numerically find to be optimal, is positive over the entire space. When computing criteria (I,II) and PPT criterion we find that they are violated for any $\gamma,\epsilon>0$ (See SM Sec. VIII [21]). The value for the criteria is shown in Fig. 3(b). For the cat states considered, our criteria (I,II) are more powerful in detecting entanglement than the entanglement criterion discussed in Refs. [11, 12], since the latter has the same entanglement bound as criterion (III), which is useful but not tight for $\rho(\psi^{+})$ (See SM Sec. VIII [21]). Also in practical considerations, different from PPT criterion and entanglement criterion based on quantum Fisher information [40] where full information of the density matrix is required, the criteria (I,II) only need partial information (a slice of Wigner function) determined by the transformation $\left(x^{\prime},p^{\prime}\right)$. Additionally, for criterion (II) the required integral region can be further narrowed down to a limited region, for example, the yellow circles in Fig. 3(a).
To bridge the gap between our theoretical criteria and practical implementations with current technology, we provide a comprehensive error analysis from two perspectives in SM Sec. IX [21]. First, we examine the impact of grid spacing when using discrete-point measurements. When implementing our criteria experimentally, the continuous Wigner function must be sampled on a discrete grid through displaced parity measurements. We demonstrate that this discretization introduces minimal errors–even for coarse sampling with grid spacing $\delta=0.5$–confirming the robustness of our approach. Second, we present a measurement strategy specifically designed for our criteria. We introduce a measurement protocol that evaluates our criteria through scanning measurements uniformly distributed over the relevant phase-space region (for example, using Fermat spiral patterns [41] or randomized measurement). This approach is statistically equivalent to single-point displaced parity measurements, thereby significantly reducing data requirements. For the paradigmatic two-mode squeezed vacuum state, our method requires two orders of magnitude fewer measurements than the four-point Bell inequality test [8, 9] to achieve comparable confidence levels.
Summary and outlook.– We have introduced criteria to detect CV bipartite entanglement from the Wigner function. By comparing them with known criteria from the literature [1, 2], we have shown cases where they are necessary and sufficient. Crucially, our criteria require only partial information on the system’s state and can be implemented experimentally by direct Wigner function measurements. Importantly, the amount of information required is no more than a single-mode Wigner tomography.
Beyond the specific examples demonstrated, each criterion targets distinct classes of entangled states. Criteria (I) and (II) provide lower bounds on the entanglement negativity and the degree of CCNR criterion violation, respectively. Criterion (III), in the meanwhile, specifically detects entangled states that exhibit Wigner negativity.
It would be interesting to investigate whether criteria similar to (I,II,III) can be found for discrete variable systems, using the Wigner function for spin states [42]. This would allow for alternative methods for the detection of entanglement between e.g. atomic ensembles [43, 44, 45, 46, 47]. We note that our approach readily extends beyond beamsplitters to encompass more general two-mode transformations of phase-space coordinates, including nonlinear operations such as two-mode squeezing.
Methods
The key insight behind criterion (I) is that if the initial state is separable, then applying partial transposition followed by a beam splitter operation should yield a physically valid state. The resulting Wigner function must satisfy the upper bound of any physical Wigner function. When this bound is violated, it reveals that our separability assumption was incorrect.
@@ Methods
Proof of criterion (I) – Consider an arbitrary transformation $(x^{\prime},p^{\prime})$ which can be determined by the real parameters $a,b,c,d,x_{0},p_{0}$. Based on these parameters, we then construct a new single-mode transformation $h:\left(\begin{array}[]{c}x\\ p\end{array}\right)\rightarrow\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\left(\begin{array}[]{c}x+\frac{X}{\cos\theta}\\ p+\frac{P}{\cos\theta}\end{array}\right)+\left(\begin{array}[]{c}x_{0}\sin\theta\\ p_{0}\sin\theta\end{array}\right)$ where $X,P$ are arbitrary constant coordinates. Then according to Eq. (9) we obtain
$W_{h(B)}^{\lambda}\left(x\sin\theta-\frac{X}{\cos\theta},p\sin\theta-\frac{P}{\cos\theta}\right)$ (14)
$=$ $W_{B}^{\lambda}\left(\left(ax+bp+x_{0}\right)\sin\theta,\left(cx+dp+p_{0}\right)\sin\theta\right)$
$=$ $W_{B}^{\lambda}\left(x^{\prime}\sin\theta,p^{\prime}\sin\theta\right).$
@@
Assume $\rho_{AB}$ is a separable state with a Wigner function given by Eq. (2). The left-hand side of criterion (I) can be rewritten as
$\iint_{-\infty}^{\infty}\sum_{\lambda}p(\lambda)W_{A}^{\lambda}(x\cos\theta,p\cos\theta)$ (15)
$\times W_{h\left(B\right)}^{\lambda}\left(x\sin\theta-\frac{X}{\cos\theta},p\sin\theta-\frac{P}{\cos\theta}\right)\text{d}x\text{d}p$
$=$ $\iint_{-\infty}^{\infty}\sum_{\lambda}p(\lambda)W_{A}^{\lambda}\left(x\cos\theta+X\sin\theta,p\cos\theta+P\sin\theta\right)$
$\times W_{h\left(B\right)}^{\lambda}\left(x\sin\theta-X\cos\theta,p\sin\theta-P\cos\theta\right)\text{d}x\text{d}p\;,$
@@
where in the last row we substituted $x\rightarrow x+X\tan\theta$ and similarly for $p$. This can be interpreted as follows: Perform a beam-splitter transformation $\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{array}\right)$ to the two modes $A$ and $h(B)$, and denote the resulting modes as $A^{\prime}:(x,p)$ and $B^{\prime}:(X,P)$. Notice that Eq. (15) is exactly the Wigner function of mode $B^{\prime}$, with mode $A^{\prime}$ integrated out. Therefore Eq. (15) should be upper-bounded by $1/(2\pi)$ [48]. Detailed proof can be found in Sec. IV of SM [21]. $\Box$
Criterion (II) exploits the fact that for separable states, the amount of correlation (with or without partial transposition) is fundamentally limited by the purities of the reduced states through the Cauchy-Schwarz inequality. When the measured correlations exceed this bound, they cannot be explained by classical correlations alone, thereby proving entanglement.
@@
$\quad\times\sqrt{\left(\iint_{R}\left[W_{f\left(\rho_{B}\right)}^{\lambda}(x\sin\theta,p\sin\theta)\right]^{2}\text{d}x\text{d}p\right)}$
@@
The physical motivation for criterion (III) coincides with fidelity calculations between two modes. For a pure product state, the fidelity between modes has a well-defined lower bound of zero. Since this expression is linear, it generalizes directly to mixed separable states. However, when the initial state is entangled, applying the same fidelity formula can yield negative values because the formula no longer represents a valid fidelity between modes for entangled states. This negative value indicates non-separability.